# A consistent, volume preserving, and adaptive mesh refinement-based framework for modeling non-isothermal gas-liquid-solid flows with phase change

### v1

@@ 2.1 PCM-gas interface tracking
Figure 1: A schematic depicting three phases: liquid (represented in blue), solid (in yellow), and gas (in white). The gray region represents the mushy region $\Omega^{\text{M}}(t)$. This is a very thin region, but it has been exaggerated for illustration purposes. A Heaviside function $H(\phi)$ is used to distinguish between the PCM $(\Omega^{\text{P}}(t)=\Omega^{\text{L}}(t)\cup\Omega^{\text{M}}(t)\cup\Omega^{\text{S}}(t))$ and gas $\Omega^{\text{G}}(t)$ regions. The liquid fraction variable $\varphi$ tracks the liquid-solid interface. The conventions adopted in this study are as follows: $H=1$ within the PCM, $H=0$ within the gas, and $10<H<$ 1 within the interface cells. $\varphi=1$ within the liquid phase, $\varphi=0$ within the solid phase, and $10<\varphi<$ 1 within the mushy region.
@@
The static computational domain $\Omega$ consists of a time varying PCM region $\Omega^{\text{P}}(t)$ and a time varying gas region $\Omega^{\text{G}}(t)$, such that $\Omega=\Omega^{\text{P}}(t)\cup\Omega^{\text{G}}(t)\;\forall t$; see Fig. 1. The PCM-gas interface $\Gamma(t)$ is tracked using the level set method. In this approach, the interface is defined as the zero contour of the level set function $\phi(\bm{\mathrm{x}},t)$
$\Gamma(t)=\{\bm{\mathrm{x}}\in\Omega\,|\,\phi(\bm{\mathrm{x}},t)=0\}.$ (1)
@@
The level set function $\phi(\mathbf{x},t)$ is taken to be a signed distance function, and defined as
$\phi(\bm{\mathrm{x}},t)=\begin{cases}>0,&\text{PCM},\\ <0,&\text{gas},\\ =0,&\text{PCM-gas interface}.\end{cases}$ (2)
@@
$\phi$ is advected with a non-divergence free velocity $\bm{\mathrm{u}}(\bm{\mathrm{x}},t)$ according to
$\frac{{\rm D}\phi}{{\rm D}t}=\frac{\partial\phi}{\partial t}+\bm{\mathrm{u}}\cdot\nabla{\phi}=0.$ (3)
Under the action of advection, $\phi$ looses its the signed distance property. The signed distance property is lost at the continuous level (due to advection), and it should not be viewed as a purely numerical problem [25]. Therefore, a reinitialization procedure [26] is employed to retain the signed distance property of $\phi(\bm{\mathrm{x}},t)$. In this work, the level set field is reinitialized at each time step. A Heaviside function $H(\bm{\mathrm{x}},t)$ is defined using the level set function $\phi(\bm{\mathrm{x}},t)$ to distinguish the PCM and gas regions. We use a smooth Heaviside function to smoothly transition material properties from the gas to the PCM domain. Its functional form reads as
@@
$H$ =$=\begin{cases}0,&\phi(\bm{\mathrm{x}})<-n_{\text{cells}}\,\Delta,\\ \frac{1}{2}\left[1+\frac{1}{n_{\text{cells}}\,\Delta}\;\phi(\bm{\mathrm{x}})+\frac{1}{\pi}\text{sin}\left(\frac{\pi}{n_{\text{cells}}\,\Delta}\;\phi(\bm{\mathrm{x}})\right)\right],&|\phi(\bm{\mathrm{x}})|\leq n_{\text{cells}}\,\Delta,\\ 1,&\text{otherwise}.\end{cases}$ (4)
@@ 2.2 Enthalpy equation
The gas and PCM domains are distinguished by the $\phi(\bm{\mathrm{x}},t)=0$ or $H(\bm{\mathrm{x}},t)=0.5$ contours. Eq. (6) introduces two key temperatures: liquidus temperature $T^{\rm liq}$ where solidification begins and solidus temperature $T^{\rm sol}$ where it completes. For pure materials, there is a single, well-defined temperature $T_{m}$ at which phase change occurs. However, alloys and glassy materials exhibit a more gradual phase change process that happens over a range of temperature $\Delta T$ = $T^{\rm liq}-T^{\rm sol}$. The enthalpy method [1] employed in this work takes the latter approach and models the melting/solidification process over an extended temperature range. The solidus and liquidus enthalpies are defined as $h^{\rm sol}=C^{\rm S}(T^{\rm sol}-T_{r})$ and $h^{\rm liq}=\bar{C}(T^{\rm liq}-T^{\rm sol})+h^{\rm sol}+L$, respectively. Here, $\bar{C}=\frac{C^{\rm S}+C^{\rm L}}{2}$ is the specific heat of the mushy region. It is taken as the average of liquid $C^{\rm L}$ and solid $C^{\rm S}$ specific heats. The latent heat of melting/fusion is denoted $L$, and the specific heat of gas is denoted $C^{\rm G}$.
@@ 2.3 Low Mach Navier-Stokes equations
Here, $\bm{\mathrm{u}}(\bm{\mathrm{x}},t)$ denotes the Eulerian velocity, $p(\bm{\mathrm{x}},t)$ represents the Eulerian pressure, and $\rho(x,t)$ and $\mu(\bm{\mathrm{x}},t)$ are spatially and temporally varying density and viscosity fields, respectively. $\rho(\bm{\mathrm{x}},t)\,\bm{\mathrm{g}}$ in the momentum Eq. (14a) represents the gravitational body force, with $\bm{\mathrm{g}}$ denoting the acceleration due to gravity. The drag force term $-A_{d}\bm{\mathrm{u}}$ is used to retard any flow in the solid region. $\bm{\mathrm{f}}_{\rm st}(\bm{\mathrm{x}},t)$ represents the surface tension force at the gas-liquid interface, which will be discussed later. The continuity Eq. (14b) provides a constraint on the velocity field and is directly derived from the conservation of mass equation:
@@ 2.5 Surface tension force
$\bm{\mathrm{f}}_{\rm st}$ $=\sigma(T)\mathcal{C}(\phi)\delta\bm{\mathrm{n}}+\nabla_{||}\sigma(T)\,\delta,$ (21a)
$\hookrightarrow\bm{\mathrm{f}}_{\rm st}$ $=\sigma(T)\mathcal{C}(\phi)\delta\bm{\mathrm{n}}+\left[\nabla{\sigma(T)}-\left(\nabla{\sigma(T)}\cdot\bm{\mathrm{n}}\right)\bm{\mathrm{n}}\right]\,\delta.$ (21b)
Here, $\mathcal{C}(\phi)=-\nabla\cdot\mbox{}\bm{\mathrm{n}}=-\nabla\cdot\mbox{}\left(\frac{\nabla{\phi}}{|\nabla{\phi}|}\right)$ is the curvature of the interface computed from the level set function $\phi$, and $\delta$ is the smoothed/mollified delta function. Implementing Eq. (21b) as is leads to spurious/parasitic velocity currents near the interface. Francois et al. [28] proposed a well-balanced scheme in which $\delta$ and $\delta\,\bm{\mathrm{n}}$ are obtained from the discrete gradient of a mollified Heaviside function $\widetilde{B}$ as $\delta=|\nabla{\widetilde{B}}|$ and $\delta\,\bm{\mathrm{n}}=\nabla\widetilde{B}$. Furthermore, the spatial gradient of the surface tension coefficient can be expressed as $\nabla{\sigma}=\frac{\rm d\sigma}{\rm dT}\,\nabla{T}$. With these substitutions, Eq. (21b) becomes
$\bm{\mathrm{f}}_{\rm st}$ $=\sigma(T)\mathcal{C}(\phi)\nabla{\widetilde{B}}+\nabla{\sigma(T)}|\nabla{\widetilde{B}}|-\left(\nabla{\sigma(T)}\cdot\frac{\nabla{\phi}}{|\nabla{\phi}|}\right)\,\nabla{\widetilde{B}},$ (22a)
$\hookrightarrow\bm{\mathrm{f}}_{\rm st}$ $=\sigma(T)\mathcal{C}(\phi)\nabla{\widetilde{B}}+\frac{\rm d\sigma}{\rm dT}\nabla{T}|\nabla{\widetilde{B}}|-\left(\nabla{T}\cdot\frac{\nabla{\phi}}{|\nabla{\phi}|}\right)\,\frac{\rm d\sigma}{\rm dT}\nabla{\widetilde{B}}.$ (22b)
@@
Figure 3: Illustration of (A) liquid fraction ($\varphi$) and extended liquid fraction ($\widetilde{\varphi}$) variables at a representative time for the case of gas bubble trapping during metal solidification considered in Sec. 7.4.
The form of the surface tension force written in Eqs. (22) is appropriate for a two-phase flow (liquid and gas) model. In three-phase flows involving liquid-solid-gas interfaces, the surface tension force form needs to be adjusted. For simulating metal manufacturing applications, gas-solid and liquid-solid surface tensions can be ignored. With this assumption, the overall three-phase dynamic can be fairly well captured, except for achieving the correct contact angle at the material triple points. In highly non-equilibrium situations like welding and metal AM, dynamic contact angle conditions are rarely known, and not imposing one is a reasonable compromise. In light of these assumptions, our three-phase surface tension force model ignores contributions from solid-gas and solid-liquid interfaces. One approach to achieving this is to multiply the two-phase surface tension force $\bm{\mathrm{f}}_{\rm st}$ by the liquid fraction variable $\varphi$, since $\varphi=0$ within the solid domain. Therefore, a possible form of the three-phase flow surface tension force reads as
@@
Although the above form of $\bm{\mathrm{f}}_{\rm st}^{\rm 3-phase}$ appears straightforward at first glance, the continuum surface tension model of Brackbill et al. introduces an additional complexity—the surface tension force acts on both sides of the liquid-gas interface. Within the gas region $\Omega^{\text{G}}$, $\varphi$ is not defined as it is not needed. In our simulations, we set $\varphi=0$ inside $\Omega^{\text{G}}$, which is an arbitrary choice. If $\varphi=0\;\forall\;\bm{\mathrm{x}}\in\Omega^{\text{G}}$ is used to define $\bm{\mathrm{f}}_{\rm st}^{\rm 3-phase}$, then the diffuse form of surface tension force acts only within one-half of the interfacial region on the PCM side. This is clearly incorrect. To resolve this situation $\varphi$ needs to be extended into the gas region. The need for $\varphi$ extension can be appreciated by considering a three-phase flow situation depicted in Fig. 2 (A). Here, a gas bubble is partially trapped in solid metal. The liquid fraction is 1 in the liquid phase and 0 in the solid phase. The liquid fraction value in the gas region is set to zero. For the three-phase surface tension model, the liquid fraction in the diffuse region on the gas side should have the same value as in the surrounding PCM region, as depicted in Fig. 2 (B). For example, on the bubble-PCM interface, $\varphi$ should equal 1 in the upper half but 0 in the lower half. It is difficult to manually set the liquid fraction within the gas region when interfaces evolve dynamically during simulation. An approach to address this issue is to solve a hyperbolic equation of the form
@@
to achieve this extension. Here, $\widetilde{\varphi}$ is the extended/extrapolated liquid fraction variable and $\bm{\mathrm{n}}=-\frac{\nabla{\phi}}{\left|\nabla{\phi}\right|}$ denotes the interfacial normal that points outward from the PCM domain and into the gas phase. This equation extrapolates/extends the liquid fraction $\varphi$ from the PCM to the gas domain along the normal direction $\bm{\mathrm{n}}$. The factor $(1-H)$ appearing in Eq. (23) ensures that the extrapolation occurs only one-way, i.e, from PCM to gas. To ensure $\widetilde{\varphi}$ is defined adequately within the smeared region, we extend it up to 5 grid cells within the gas domain. This requires integrating Eq. (23) for approximately 15 pseudo times steps of size $\Delta\tau=0.3\Delta$. Here, $\Delta$ denotes the cell size and a CFL number of 0.3 is employed to ensure convective stability. An example of original and extrapolated liquid fraction variables, $\varphi$ and $\widetilde{\varphi}$, respectively, during a three-phase flow simulation considered in Sec. 7.4 is shown in Fig. 2. We remark that $\widetilde{\varphi}$ is used only to apply surface tension forces along the gas-liquid interface. Wherever else it is applicable, for instance, when defining thermophysical properties (using Eq. (13b)), the original liquid fraction variable $\varphi$ is used. In light of this discussion, the form of the three-phase surface tension force used in this work reads as

### v2

@@ 2 Mathematical formulation
In this section, we succinctly write the low Mach enthalpy equations derived in our previous work [1] for reader’s convenience. We also present a thermodynamic justification for the low Mach enthalpy model, a three-phase surface tension model, and a heat source term in the enthalpy equation, which were not discussed in our previous work [1].
@@ 2.1 PCM-gas interface tracking
Figure 1: A schematic depicting three phases: liquid (represented in blue), solid (in yellow), and gas (in white). The gray region represents the mushy region $\Omega^{\text{M}}(t)$. This is a very thin region, but it has been exaggerated for illustration purposes. A Heaviside function $H({\color[rgb]{0,0,0}\Phi})$ is used to distinguish between the PCM $(\Omega^{\text{P}}(t)=\Omega^{\text{L}}(t)\cup\Omega^{\text{M}}(t)\cup\Omega^{\text{S}}(t))$ and gas $\Omega^{\text{G}}(t)$ regions. The liquid fraction variable $\varphi$ tracks the liquid-solid interface. The conventions adopted in this study are as follows: $H=1$ within the PCM, $H=0$ within the gas, and $10<H<$ 1 within the interface cells. $\varphi=1$ within the liquid phase, $\varphi=0$ within the solid phase, and $10<\varphi<$ 1 within the mushy region.
@@
The static computational domain $\Omega$ consists of a time varying PCM region $\Omega^{\text{P}}(t)$ and a time varying gas region $\Omega^{\text{G}}(t)$, such that $\Omega=\Omega^{\text{P}}(t)\cup\Omega^{\text{G}}(t)\;\forall t$; see Fig. 1. The PCM-gas interface $\Gamma(t)$ is tracked using the level set method. In this approach, the interface is defined as the zero contour of the level set function ${\color[rgb]{0,0,0}\Phi}(\bm{\mathrm{x}},t)$
$\Gamma(t)=\{\bm{\mathrm{x}}\in\Omega\,|\,{\color[rgb]{0,0,0}\Phi}(\bm{\mathrm{x}},t)=0\}.$ (1)
@@
The level set function ${\color[rgb]{0,0,0}\Phi}(\mathbf{x},t)$ is taken to be a signed distance function, and defined as
${\color[rgb]{0,0,0}\Phi}(\bm{\mathrm{x}},t)=\begin{cases}>0,&\text{PCM},\\ <0,&\text{gas},\\ =0,&\text{PCM-gas interface}.\end{cases}$ (2)
@@
${\color[rgb]{0,0,0}\Phi}$ is advected with a non-divergence free velocity $\bm{\mathrm{u}}(\bm{\mathrm{x}},t)$ according to
$\frac{{\rm D}{\color[rgb]{0,0,0}\Phi}}{{\rm D}t}=\frac{\partial{\color[rgb]{0,0,0}\Phi}}{\partial t}+\bm{\mathrm{u}}\cdot\nabla{{\color[rgb]{0,0,0}\Phi}}=0.$ (3)
Under the action of advection, ${\color[rgb]{0,0,0}\Phi}$ looses its the signed distance property. The signed distance property is lost at the continuous level (due to advection), and it should not be viewed as a purely numerical problem [25]. A reinitialization procedure proposed by Sussman et al. [26] is employed to retain the signed distance property of ${\color[rgb]{0,0,0}\Phi}(\bm{\mathrm{x}},t)$. This is achieved by computing the steady-state solution to the Hamilton-Jacobi equation
$\frac{\partial\Phi}{\partial\tau}+\textrm{sgn}\left(\widetilde{\Phi}\right)\left(\|\nabla\Phi\|-1\right)=0,$ (4a)
$\Phi(\bm{\mathrm{x}},\tau=0)=\widetilde{\Phi}(\bm{\mathrm{x}}).$ (4b)
At the end of a physical time step, Eq. (4a) is evolved in pseudo-time $\tau$, which, at steady state, produces a signed distance function satisfying the Eikonal equation $\|\nabla\Phi\|=1$. Here, sgn denotes the sign of $\widetilde{\Phi}$, which is either $1$, $-1$, or 00. To mitigate mass loss, we employ the subcell-fix method proposed by Min [27] and enforce an immobility condition near the zero level set as described by Son [28]. For more details about the reinitialization procedure used in this work, please refer to our previous papers [29, 30]. The level set field is reinitialized at each time step. A Heaviside function $H(\bm{\mathrm{x}},t)$ is defined using the level set function ${\color[rgb]{0,0,0}\Phi}(\bm{\mathrm{x}},t)$ to distinguish the PCM and gas regions. We use a smooth Heaviside function to smoothly transition material properties from the gas to the PCM domain. Its functional form reads as
@@
$H$ =$=\begin{cases}0,&{\color[rgb]{0,0,0}\Phi}(\bm{\mathrm{x}})<-n_{\text{cells}}\,\Delta,\\ \frac{1}{2}\left[1+\frac{1}{n_{\text{cells}}\,\Delta}\;{\color[rgb]{0,0,0}\Phi}(\bm{\mathrm{x}})+\frac{1}{\pi}\text{sin}\left(\frac{\pi}{n_{\text{cells}}\,\Delta}\;{\color[rgb]{0,0,0}\Phi}(\bm{\mathrm{x}})\right)\right],&|{\color[rgb]{0,0,0}\Phi}(\bm{\mathrm{x}})|\leq n_{\text{cells}}\,\Delta,\\ 1,&\text{otherwise}.\end{cases}$ (5)
@@ 2.2 Enthalpy equation
Here, $C^{\rm L}$, $C^{\rm S}$ and $C^{\rm G}$ are the specific heats of liquid, solid and gas, respectively. The gas and PCM domains are distinguished by the ${\color[rgb]{0,0,0}\Phi}(\bm{\mathrm{x}},t)=0$ or $H(\bm{\mathrm{x}},t)=0.5$ contours. Eq. (7) introduces two key temperatures: liquidus temperature $T^{\rm liq}$ where solidification begins and solidus temperature $T^{\rm sol}$ where it completes. For pure materials, there is a single, well-defined temperature $T_{m}$ at which phase change occurs. However, alloys and glassy materials exhibit a more gradual phase change process that happens over a range of temperature $\Delta T$ = $T^{\rm liq}-T^{\rm sol}$. The enthalpy method [1] employed in this work takes the latter approach and models the melting/solidification process over an extended temperature range. The solidus and liquidus enthalpies are defined as $h^{\rm sol}=C^{\rm S}(T^{\rm sol}-T_{r})$ and $h^{\rm liq}=\bar{C}(T^{\rm liq}-T^{\rm sol})+h^{\rm sol}+L$, respectively. Here, $\bar{C}=\frac{C^{\rm S}+C^{\rm L}}{2}$ is the specific heat of the mushy region. It is taken as the average of liquid $C^{\rm L}$ and solid $C^{\rm S}$ specific heats. The latent heat of melting/fusion is denoted $L$, and the specific heat of gas is denoted $C^{\rm G}$.
@@ 2.3 Low Mach Navier-Stokes equations
Here, $\bm{\mathrm{u}}(\bm{\mathrm{x}},t)$ denotes the Eulerian velocity, $p(\bm{\mathrm{x}},t)$ represents the Eulerian pressure, and $\rho(x,t)$ and $\mu(\bm{\mathrm{x}},t)$ are spatially and temporally varying density and viscosity fields, respectively. $\rho(\bm{\mathrm{x}},t)\,\bm{\mathrm{g}}$ in the momentum Eq. (15a) represents the gravitational body force, with $\bm{\mathrm{g}}$ denoting the acceleration due to gravity. The drag force term $-A_{d}\bm{\mathrm{u}}$, commonly referred to as the Carman-Kozeny force [11] is used to retard any flow in the solid region. $\bm{\mathrm{f}}_{\rm st}(\bm{\mathrm{x}},t)$ represents the surface tension force at the gas-liquid interface, which will be discussed later. The continuity Eq. (15b) provides a constraint on the velocity field and is directly derived from the conservation of mass equation:
@@ 2.5 Surface tension force
$\bm{\mathrm{f}}_{\rm st}$ $=\sigma(T)\mathcal{C}({\color[rgb]{0,0,0}\Phi})\delta\bm{\mathrm{n}}+\nabla_{||}\sigma(T)\,\delta,$ (22a)
$\hookrightarrow\bm{\mathrm{f}}_{\rm st}$ $=\sigma(T)\mathcal{C}({\color[rgb]{0,0,0}\Phi})\delta\bm{\mathrm{n}}+\left[\nabla{\sigma(T)}-\left(\nabla{\sigma(T)}\cdot\bm{\mathrm{n}}\right)\bm{\mathrm{n}}\right]\,\delta.$ (22b)
Here, $\mathcal{C}({\color[rgb]{0,0,0}\Phi})=-\nabla\cdot\mbox{}\bm{\mathrm{n}}=-\nabla\cdot\mbox{}\left(\frac{\nabla{{\color[rgb]{0,0,0}\Phi}}}{|\nabla{{\color[rgb]{0,0,0}\Phi}}|}\right)$ is the curvature of the interface computed from the level set function ${\color[rgb]{0,0,0}\Phi}$, and $\delta$ is the smoothed/mollified delta function. Implementing Eq. (22b) as is leads to spurious/parasitic velocity currents near the interface. Francois et al. [32] proposed a well-balanced scheme in which $\delta$ and $\delta\,\bm{\mathrm{n}}$ are obtained from the discrete gradient of a mollified Heaviside function $\widetilde{B}$ as $\delta=|\nabla{\widetilde{B}}|$ and $\delta\,\bm{\mathrm{n}}=\nabla\widetilde{B}$. We use Peskin’s four-point regularized delta function [33] to mollify the Heaviside function as
$\widetilde{B}(\bm{\mathrm{x}},t)=\int_{\Omega}H(\bm{\mathrm{y}},t)\delta(\bm{\mathrm{y}}-\bm{\mathrm{x}}){\rm{d}}\bm{\mathrm{y}},$ (23)
in which $\delta(\bm{\mathrm{y}}-\bm{\mathrm{x}})$ represents the four-point regularized delta function.
Furthermore, the spatial gradient of the surface tension coefficient can be expressed as $\nabla{\sigma}=\frac{\rm d\sigma}{\rm dT}\,\nabla{T}$. With these substitutions, Eq. (22b) becomes
$\bm{\mathrm{f}}_{\rm st}$ $=\sigma(T)\mathcal{C}({\color[rgb]{0,0,0}\Phi})\nabla{\widetilde{B}}+\nabla{\sigma(T)}|\nabla{\widetilde{B}}|-\left(\nabla{\sigma(T)}\cdot\frac{\nabla{{\color[rgb]{0,0,0}\Phi}}}{|\nabla{{\color[rgb]{0,0,0}\Phi}}|}\right)\,\nabla{\widetilde{B}},$ (24a)
$\hookrightarrow\bm{\mathrm{f}}_{\rm st}$ $=\sigma(T)\mathcal{C}({\color[rgb]{0,0,0}\Phi})\nabla{\widetilde{B}}+\frac{\rm d\sigma}{\rm dT}\nabla{T}|\nabla{\widetilde{B}}|-\left(\nabla{T}\cdot\frac{\nabla{{\color[rgb]{0,0,0}\Phi}}}{|\nabla{{\color[rgb]{0,0,0}\Phi}}|}\right)\,\frac{\rm d\sigma}{\rm dT}\nabla{\widetilde{B}}.$ (24b)
@@
Figure 3: Illustration of (A) liquid fraction ($\varphi$) and (B)  extended liquid fraction ($\widetilde{\varphi}$) variables at a representative time for the case of gas bubble trapping during metal solidification considered in Sec. 7.4.
The form of the surface tension force written in Eqs. (24) is appropriate for a two-phase flow (liquid and gas) model. In three-phase flows involving liquid-solid-gas interfaces, the surface tension force form needs to be adjusted. For simulating metal manufacturing applications, gas-solid and liquid-solid surface tensions can be ignored. With this assumption, the overall three-phase dynamic can be fairly well captured, except for achieving the correct contact angle at the material triple points. In highly non-equilibrium situations like welding and metal AM, dynamic contact angle conditions are rarely known, and not imposing one is a reasonable compromise. In light of these assumptions, our three-phase surface tension force model ignores contributions from solid-gas and solid-liquid interfaces. One approach to achieving this (ignoring surface tension of the solid phase) is to multiply the two-phase surface tension force $\bm{\mathrm{f}}_{\rm st}$ by the liquid fraction variable $\varphi$, since $\varphi=0$ within the solid domain. Therefore, a possible form of the three-phase flow surface tension force reads as
@@
Although the above form of $\bm{\mathrm{f}}_{\rm st}^{\rm 3-phase}$ appears straightforward at first glance, the continuum surface tension model of Brackbill et al. introduces an additional complexity—the surface tension force acts on both sides of the liquid-gas interface. Within the gas region $\Omega^{\text{G}}$, $\varphi$ is not defined as it is not needed. In our simulations, we set $\varphi=0$ inside $\Omega^{\text{G}}$, which is an arbitrary choice. If $\varphi=0\;\forall\;\bm{\mathrm{x}}\in\Omega^{\text{G}}$ is used to define $\bm{\mathrm{f}}_{\rm st}^{\rm 3-phase}$, then the diffuse form of surface tension force acts only within one-half of the interfacial region on the PCM side. This is clearly incorrect. To resolve this situation $\varphi$ needs to be extended into the gas region. The need for $\varphi$ extension can be appreciated by considering a three-phase flow situation depicted in Fig. 2 (A). Here, a gas bubble is partially trapped in solid metal. The liquid fraction is 1 in the liquid phase and 0 in the solid phase. The liquid fraction value in the gas region is set to zero. For the three-phase surface tension model, the liquid fraction in the diffuse region on the gas side should have the same value as in the surrounding PCM region, as depicted in Fig. 2 (B). For example, on the bubble-PCM interface, $\varphi$ should equal 1 in the upper half but 0 in the lower half. It is difficult to manually set the liquid fraction within the gas region when interfaces evolve dynamically during simulation. An approach to address this issue is to solve a hyperbolic equation of the form [34]
@@
to achieve this extension. Here, $\widetilde{\varphi}$ is the extended/extrapolated liquid fraction variable and $\bm{\mathrm{n}}=-\frac{\nabla{{\color[rgb]{0,0,0}\Phi}}}{\left|\nabla{{\color[rgb]{0,0,0}\Phi}}\right|}$ denotes the interfacial normal that points outward from the PCM domain and into the gas phase. This equation extrapolates/extends the liquid fraction $\varphi$ from the PCM to the gas domain along the normal direction $\bm{\mathrm{n}}$. The factor $(1-H)$ appearing in Eq. (25) ensures that the extrapolation occurs only one-way, i.e, from PCM to gas. To ensure $\widetilde{\varphi}$ is defined adequately within the smeared region, we extend it up to 5 grid cells within the gas domain. This requires integrating Eq. (25) for approximately 15 pseudo times steps of size $\Delta\tau=0.3\Delta$. Here, $\Delta$ denotes the cell size and a CFL number of 0.3 is employed to ensure convective stability. An example of original and extrapolated liquid fraction variables, $\varphi$ and $\widetilde{\varphi}$, respectively, during a three-phase flow simulation considered in Sec. 7.4 is shown in Fig. 2. We remark that $\widetilde{\varphi}$ is used only to apply surface tension forces along the gas-liquid interface. Wherever else it is applicable, for instance, when defining thermophysical properties (using Eq. (14b)), the original liquid fraction variable $\varphi$ is used. In light of this discussion, the form of the three-phase surface tension force used in this work reads as